OpenStudy (anonymous):

create your own polynomial with a degree of 3 that when divided by x+2 has a remainder of -4

OpenStudy (loser66):

give me a polynomial degree 1

OpenStudy (anonymous):

x^2 i think

OpenStudy (loser66):

it's degree2 friend, degree is exponent, you give x^2 is not degree 1

OpenStudy (anonymous):

k

OpenStudy (loser66):

you have (x+2) is a polynomial degree1, give me 2 more

OpenStudy (anonymous):

thats all i have

OpenStudy (loser66):

ok, i take your x^2, now x^2 (x+2) = -4 open the LHS , you have x^3 +2x^2 =-4 plus 4 both sides x^3 +2x^2 +4 =0 that 's what you need

OpenStudy (anonymous):

okay